OpenStudy (anonymous):

12x Squared over 2x-3 divided by 3x-12 over 8x-12

OpenStudy (anonymous):

can you put this in equation form using the button below?

OpenStudy (anonymous):

I tried Its Not Typing It in Like its on My Paper

OpenStudy (anonymous):

hmm then is it possible for you to draw it?

OpenStudy (anonymous):

ill try here it go

OpenStudy (anonymous):

\[\frac{12x^2 }{ (2x-3)(3x-12)(8x-12) }\]

OpenStudy (anonymous):

does it look like that?

OpenStudy (anonymous):

no the 12x squared is over the 2x-3

OpenStudy (anonymous):

and the 3x-12 is over the 8x-12

OpenStudy (anonymous):

\[\frac{ 12x^2 }{ 2x-3 } \div \frac{ 3x-12 }{ 8x-12 }\]

OpenStudy (anonymous):

like that?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

okay first can you simplify anything?

OpenStudy (anonymous):

I dont know i have only did questions like this twice before on a quiz and i always get them wrong

OpenStudy (anonymous):

okay lets look at 8x-12, do you know a number that multiplies to both 8 and 12

OpenStudy (anonymous):

no 4 because 8 divided by 4 is 2 and 12 divided by 4 is 3

OpenStudy (anonymous):

okay that's awesome so if you take out the four what do you get?

OpenStudy (anonymous):

2 and 3

OpenStudy (anonymous):

write that in equation form. and try to write this down on a paper so you know what your doing and can follow it easier

OpenStudy (anonymous):

so basically 8x-12 becomes (4)(2x-3)

OpenStudy (anonymous):

so it would be 12x^2 /2 and 3x-12/3??

OpenStudy (anonymous):

ummm no. so far all we have done is make your equation look like this -->\[\frac{ 12x^2 }{ 2x-3 } \div \frac{ 3x-12 }{ (4)(2x-3)}\]

OpenStudy (anonymous):

oh okay

OpenStudy (anonymous):

did you know